OpenStudy (anonymous):

compute the exact value of sin(tan^-1(1/3))

OpenStudy (sasogeek):

use ur calculator :P

OpenStudy (anonymous):

well..i got 0.316227766

OpenStudy (zarkon):

\[\frac{\sqrt{10}}{10}\]

OpenStudy (anonymous):

i lost my calculator :(

OpenStudy (zarkon):

you don't need a calculator

OpenStudy (anonymous):

0_o

OpenStudy (zarkon):

let \[\theta=\tan^{-1}(1/3)\] then\[\tan(\theta)=1/3\] draw a right triangle with the above numbers then evaluate \(\sin(\theta)\)